OpenStudy (anonymous):

-7 + 4 is equivalent to? -11 3 -3 11

OpenStudy (anonymous):

suppose you owe someone 7 dollars and you pay them 4 dollars how much do you owe them now?

OpenStudy (anonymous):

4-7=?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3 or -3

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

@jack63 you are corrrrecttt

OpenStudy (anonymous):

thanks

mathslover (mathslover):

@jack63 whenever you are told problem like -4 + 7 then write them as 7 - 4 which is easy : 3 . Here you have : - 7 + 4 , so write it as 4 - 7 you will get -3

OpenStudy (anonymous):

Which sum will be neither positive nor negative? 3 + 2 -4 + (-4) -6 + 6 7 + (-6)

mathslover (mathslover):

Can you name the no. which is neither positive nor negative @jack63

OpenStudy (anonymous):

um no

mathslover (mathslover):

Is -1 negative or positive?

OpenStudy (anonymous):

negative

mathslover (mathslover):

Great! Is 1 positive or negative?

OpenStudy (anonymous):

postitive

OpenStudy (anonymous):

-3

mathslover (mathslover):

Ok! Now. is 0 negative or positive?

OpenStudy (anonymous):

ohh neither

mathslover (mathslover):

That is it. now chose from the options which results into 0 that is neither positive or negative.

OpenStudy (anonymous):

the second one

mathslover (mathslover):

-4 + (-4 ) = -4 - 4 ... now?

mathslover (mathslover):

Note that ... a + (-b) means a -b

OpenStudy (anonymous):

okey

mathslover (mathslover):

Now try it again.

OpenStudy (anonymous):

the third one

mathslover (mathslover):

Excellent work! You're right!

OpenStudy (anonymous):

What is the sum of -45 and 17? -28 -62 28 -32

mathslover (mathslover):

Did you try it first ?

OpenStudy (anonymous):

ooooooo -28 sooooooooooooooooooooooorrrrrrrrrrrrrrrrrrrrrrrrry

mathslover (mathslover):

Right! Good.

mathslover (mathslover):

Do some practice on these type of questions. best of luck!